S Band Narrow Gregorian Dome Receiver Log

as of Monday, 29-Nov-2021 16:26:46 AST

Go to Most Recent Entry

ANS anolla@naic.edu 10/7/99 Problem: There it was a helium leak in the cryo system. Correction: The leak was in the line filter, so filter was replaced.
ANS anolla@naic.edu 1/11/2000 Problem: High Tsys. in both channels.Mike Nolan reported it. Correction: We found water condensed at the quartz domes in the dewar.We recommend to pressurize the WG.
ANS anolla@naic.edu 8/9/2000 Dewar warmed up due to work on compressor (It was off for 5 minutes),but Dewar didn't hold its vacuum and warmed up.
ANS anolla@naic.edu 10/ 10/2000 W Genter installed for SETI people the 2270 mhz filters on receiver.
ANS anolla@naic.edu 10/25/2000 W Genter installed back the 2380 mhz filters on receiver.
ANS anolla@naic.edu 1/22/2001 There is a cryo leak on line 9 and is about 12 pounds / day.
ANS anolla@naic.edu 2/27/2001 The 2270 mhz filters were installed in the receiver for SETI people.
ANS anolla@naic.edu 3/2/2001 The 2270 mhz filters removed , 2380 mhz ones back..
ANS anolla@naic.edu 3/12/2001 The cryo comp failed due to a power glitch.Jose Rosa & WG came on 3/11/2001 to cooled it down.
ANS anolla@naic.edu 3/19/2001 JR brought down the post amp chassis in order to install a transfer sw to sw between the regular RF filters and the SETI (2270 mhz) people ones.
ANS anolla@naic.edu 5/24/2001 The remote cryo temp was reading wrongly, we check and was that accidentally we left the rem/loc sw on the local pos instead remote.
ANS anolla@naic.edu 7/27/2001 Cryo comp ED 012 replaced with ED 010. ED 012 taken down for maint.
ANS anolla@naic.edu 8/30/2001 Cryo turned off in order to replace the crosshead, needs replacement after two years !!!!!!!
ANS anolla@naic.edu 9/4/2001 Crosshead replaced (ED 007) with a cleaned one (ED 015). We took crosshead ED 014 to replace the ED 007, but found that crosshead was without preassure and had loosen screws on it.
ANS anolla@naic.edu 9/10/2001 The system was cooled down on 9/6/2001, but on 9/7/2001 JR found that the crosshead (ED 015) was contaminated, so he turned it off. Then today the crosshead was removed from the Dewar.
ANS anolla@naic.edu 9/17/2001 The Dewar crosshead installed (ED 014) on 9/14/2001 and also cooled down. But on 9/15/2001 due to a power failure the compressor was tripped off, so JR came to cool it down again the same day.
ANS anolla@naic.edu 11/12/2001 Cryo comp ED 010 replaced with cryo comp ED 009.
ANS anolla@naic.edu 3/1/2002 NSB recvr filter sw to SETI people ( 2270 Mhz)
ANS anolla@naic.edu 3/4/2002 NSB recvr filter sw to normal ( 2380 Mhz)
ANS anolla@naic.edu 5/13/2002 Due to power dip system got hot due to cryocomp failure.
ANS anolla@naic.edu 5/17/2002 System was hot due to a power dip that tripped off compressor.Was cooled .
ANS anolla@naic.edu 7/31/2002 On 6/21/2002 the cryocomp 009 failed (cooling fan stop working) , it was replaced by cryo comp 011.
LSWlwray@naic.eduFriday, August 16, 2002 16:07:06Power glitch caused 3-receiver compressor to shut down, by the time it was noticed it was 1.5 hours later. The SBN warmed up to >200K by 11 am, but by 4:30 pm it had been cooled.
ansanolla@naic.eduMonday, September 9, 2002 16:00:22Due to a power dip on 9/8/2002 at 08:30 the cryo comp shut off , it was not noticed for several hours.WG had to come to cool it down.
ansanolla@naic.eduMonday, September 16, 2002 13:23:18The 16 cryo stage temp is cyclin , Ivan sw the cryopump driver to see if it helps.
ansanolla@naic.eduWednesday, September 18, 2002 12:45:10The Dewar was shut off in order to replace cryopump , cryo system was cycling.
ansanollaThursday, September 19, 2002 13:04:37The Dewar chann B amp third stage has a short on it. So Dewar is been warmed in order to check it.
ans anolla@naic.edu Tuesday, September 24, 2002 11:54:49 Dewar opened up and the wiring completely re-done. There were shorts in the old wiring, not found until the temperature had been reclycled. Also chann B (#101) rf amp was relpaced it has the 1 st & 3 rd stages shorted out ( measuring more than 30 ma ). Currently in the dewar, the amps are: Ch A: S/N 102, Ch B: S/N 103. S/N 101 to be repaired.
ansanolla@naic.eduThursday, September 26, 2002 14:15:19Dewar 16 cryo temp stage cycling between 35 & 49 k.
LSWlwray@naic.eduWednesday, October 2, 2002 13:54:40Dewar has been warmed up. Receiver still functional, but Tsys will be high.
ansanolla@naic edu.Tuesday, October 8, 2002 09:33:07The cryopump was replaced again. This time the 16 stage was lower than 35 ( 13.5)
ansanolla@naic.eduWednesday, October 9, 2002 12:47:21Dewar valve closed , it was open since monday .The cryopump was replaced , Dewar cryo temps are now ok. ( 13 & 74 )
ansanolla@naic edu.Friday, December 6, 2002 12:58:41System was set for SETI run , it was tuned to 2270 mhz.But was returned to 2380 mhz today.
ansanolla@naic edu.Monday, December 30, 2002 15:47:55There was a complain that Chann A T sys is 5k higher than Chann B. We check for water inside the waveguide quarz , but there was not water on it . Did the same check out on chann B and no water on it again.
ansanolla@naic.eduThursday, January 2, 2003 14:24:02The SB xmitter doppler was replced by the 10 mhz one (narrow).
lswlwray@naic.eduMonday, March 3, 2003 16:27:23Changed attenuators after post amp from: 0/0 dB (A/B) to 0/4 dB, rebalancing levels after upstairs RF switch change.
lswlwray@naic.eduThursday, March 6, 2003 16:34:21Cheo and Nolla noticed bronze filings in the cryopump window. pump is working ok, the temperature is holding. the filings are probably from the bushings in the motor. plan to warm up the receiver on friday and replace on monday or tuesday. Last change of cryopump was Oct 3 2002.
ansanolla@naic edu.Monday, March 10, 2003 11:47:12Cryo turn off in order to replace cryopump , is highly contaminated.
ansanolla@naic.eduThursday, March 13, 2003 14:25:35System was cooled down .Its cryo temps are 15.5 & 76 k. Its Tsys is 27 on chn A and 24 k on chn B.
lswlwray@naic.eduWednesday, March 26, 2003 16:26:31Added XB to compressor 011, now it has XB/SBN/LBN.
ansanolla@naic.eduTuesday, April 8, 2003 12:54:09System hooked up for SETI people.Noise adde3d to it thru a D.C.
ansanolla@naic.eduMonday, April 28, 2003 11:51:13The 16 stage looks to be warming up ,still on observation.
ansanolla@naic.eduTuesday, July 1, 2003 15:55:58The SBN down for modifications.
ansanolla@naic.eduMonday, July 14, 2003 10:41:55The system was shut off in order to check the Tsys of channel A is 3 higher than normal.
ansanolla@naic.eduThursday, July 17, 2003 15:56:19System been cooled down after a check out on channel A by Ganesh & Lisa.So next Monday we will check it cooled.
ansanolla@naic.eduFriday, July 18, 2003 13:41:46The SBN is back into service.Was cooled down & tis Tsys were: Chann A 27.2 K Chann B 26.9 K
ansanolla@naic.eduWednesday, July 23, 2003 09:05:43System warmed up due to power dip. Operator tried to restart the cryo comp but was not able. Later we found that a relay in the reset ckt was bad.
ansanolla@naic.eduTuesday, September 30, 2003 15:22:50The SBN shutter status indicator pilot ligth for the shutter close condition was replaced by an LED.
ansanolla@naic.eduWednesday, October 1, 2003 15:46:49The shutter xmitter interlock system is intermmitant.We found a broken wire at the termial strip of the shutter box .Also we replaced the K1 relay ,it was reading around 50 ohm across it.
ansanolla@naic.eduWednesday, October 15, 2003 12:27:20Shorts at the transition were swap in order to swap polarity on receiver , also cables at the post amp.
ansanolla@naic.eduMonday, November 10, 2003 13:17:01System warmed up on Saturday due to cryocomp failure (ED 011), it was replaced by cryocomp ED 010 (SBH one).
lswlwray@naic.eduFriday, January 16, 2004 09:52:01Wednesday 14th, in evening. short, 10's of kW blast from transmitter through open waveguide into receiver floor. unknown as to the consequences.
lswlwray@naic.eduFriday, January 30, 2004 12:47:56problem with the cryogenic lines serving the SBN and XB. There is a huge leak somewhere. The receivers are warm until further notice.
ansanolla@naic.eduTuesday, February 3, 2004 14:53:56There it was a huge leak in cryo line 13. It was replaced with a new 3/4" cryo line.System was cooled down.Back into service.
ansanolla@naic.eduWednesday, February 25, 2004 07:59:35The kevlar window on horn replaced . also the quartz windows in both channaels were widly wet , were dry out.
lswlwray@naic.eduWednesday, February 25, 2004 16:37:09Cheo and Nolla noticed water on the rf pickup probe near the crystal barrier in the waveguide. They cleaned both pickups. The kapton seal over the horn had not been sealed, probably since Feb 15 when the surveying was done.
lswlwray@naic.eduTuesday, June 1, 2004 16:01:24System had warmed up during power outage on Monday May 31 at 1am. Started cooldown today, will be cold by 8pm.
lswlwray@naic.eduWednesday, June 2, 2004 16:12:36This morning, the compressor that runs the SBN and XB failed (flappers). The XB and SBN receivers warmed up. They were put onto the SBW's compressor (the receiver is down due to cryo problems) and cooled. Both receivers are ready as of 4pm.
lswlwray@naic.eduThursday, June 3, 2004 15:58:49SBN Tx run today, receiver was cold and ready to go.
lswlwray@naic.eduThursday, June 10, 2004 08:07:32June 9: Cheo found the dewar sweating and the temperatures reading 80K on the 16K stage. VPG is not working. Since June 3 when receiver was cooled back down on a new compressor, the temperature readings have been erratic. No idea why the receiver warmed up suddenly on the 9th.
ansanolla@naic.eduFriday, July 2, 2004 11:28:03System down due to power shut down (main power line shorted at rotary joint).
ansanolla@naic.eduTuesday, July 6, 2004 15:38:19System been cooled.
ansanolla@naic.eduWednesday, July 7, 2004 14:00:45System back to normal after being down since 7/2/2004.Also there is a problem with the cryo temp 16 reading stage , the signal coming from the 16stage diode is not reaching the ckt. to read the cryotemp. The 70 stage is working properly so we decide to leave the 70 stage hooked up as our reference.
ansanolla@naic.eduTuesday, July 13, 2004 08:47:50The 16 stage diode bad inside the Dewar, so the 70 stage diode left as a common monitor until further notice.
ansanolla@naic.eduThursday, July 15, 2004 16:18:36Cryocomp swap from ED 012 to ED008 , 012 had a lot vibration on it.Been vac pumped , it lost it's vac too soon?
ansanolla@naic.eduFriday, July 23, 2004 16:03:18Cryocomp power feed was removed from shack 1 to shack 2, Dewar lost vacuum , so vacpump on it for a day.Vac valve was closed today during the morning time.
ansanolla@naic.eduTuesday, July 27, 2004 13:37:50Due to power dip cryocomp ED 008 on 7/25/2004 at 13:35 stop working .It's cooling fan quit working overheating the unit.Spare cryocomp ED 12 attached to system , but there was a leaky hose ( return one) in these cryocomp making the system to warm up again.So for these reason SBN & XB were out of service.
ansanolla@naic.eduTuesday, July 27, 2004 13:45:54System been cooled.Cryocomp ED 009 on it.
ansanolla@naic.eduFriday, July 30, 2004 11:39:31On 7/28/2004 cryopump got stuck due to contamination . Was purged & contamination ran away.
ansanolla@naic.eduThursday, September 2, 2004 12:47:24Dewar down for repairs. Bad vpg meter & 16 deg stage diode.
ansanolla@naic.eduFriday, September 17, 2004 15:09:22System back into service .Chann A Tsys is 24.9 k deg & B is 26.7 k deg.
W Genterwgenter@naic.eduSaturday, September 18, 2004 14:15:57The SBN receiver was in the cryo lab to find out the reason this receiver would warm up soon after power was removed. We first evacuated the receiver and found that the vacuum would not go below 20 uc. A possible leak was found in the VACION elbow and was sealed with torr-seal. Next we disassembled the receiver to inspect the O-Rings, replace the VPG and check the 16K Diode which was not wroking. Several O-Rings were found to be in questionable condition and we decided to replace all O-Rings since this had not been done since the original installation. The O-Rings were ordered. While waiting for the O-Rings, we inspected the Temp Diode and found that one of the leads was broken. This was repaired. When the O-Rings arrived, a new operational VPG was installed. The dewar was cleaned. Additional cold sync straps were placed on the Isolators and input cables. Ganesh tested the amplifiers and they were re-installed in the dewar. Following assembly, we evacuated the dewar and this time the vacuum was down to 2 uc without the refirgerator operating. The SBN was reinstalled in the Gregorian on September 17, 2004 for the RADAR run this weekend.
ansanolla@naic.eduMonday, September 27, 2004 11:40:07Dewar lost vacuum , so receiver warmed up.We pour vac seal into the place that Bill found leaky and it fix it.
ansanolla@naic.eduWednesday, September 29, 2004 14:34:31Cryocomp swap , it was ED 009 and was taken down for modification . Instead ED 008 was installed.
ansanolla@naic.eduMonday, October 11, 2004 12:41:24System was warmed on 10/9/2004 due to power failure.Was cooled down by BG and Ganesh.
ansanolla@naic.eduTuesday, October 12, 2004 13:05:58Dewar lost vacuum , suspect thru the VPG meter ,is loose.Meanwhile been keep under observation during the next two weeks.
ansanolla@naic.eduWednesday, October 13, 2004 11:14:05Quartz domes full of water, A worst.It's Dewar vacuum was OK.
ansanolla@naic.eduFriday, October 29, 2004 15:32:16Dewar up after been taken down on Wednesday in order to fix the VPG meter that was loose causing a vacuum leak.So system back to normal.
ansanolla@naic.eduTuesday, November 23, 2004 16:23:11There was an impedance mismatch in channel post amp that was causing a ripple thru the band.Ganesh fix it adding a 3 db pad at the input of the amp , also rf cables were made shorter.
ansanolla@naic.eduTuesday, December 28, 2004 11:34:39System off in order to replace Dewar crosshead.
ansanolla@naic.eduThursday, December 30, 2004 13:49:26Cryohead replaced , it was ED014 ,now is ED007.System back to normal.
ansanolla@naic.eduTuesday, January 4, 2005 15:54:19System warmed up due that cryohead broke it's holding screws after its replacement. Cryo head replaced with ED019.System looks contaminated but usable after cooldown.
ansanolla@naic.eduThursday, January 13, 2005 12:49:24System was hooked up back to cryocomp ED 008 , so as before XB & SBN are with this cryocomp.
lswlwray@naic.eduThursday, January 27, 2005 12:43:16After some high tsys readings in Ch A (10K above normal), we checked all RF and cal cable connections and found all was well. We checked inside the waveguide just after the cal coupler and there was water collecting on the quartz bulb around the pickup into the dewar. There was more water in Ch A than B, we wiped both dry. Checked kapton cover on receiver horn, it was intact. Afterwards tsys showed Ch A still high.
lswlwray@naic.eduThursday, January 27, 2005 15:51:36Stability testing: Nolla found loose connector on IFLO cable Ch A, underneath heatshrink jacket, fixed. A second problem: Tsys in Ch A changes from 33K to 23K in gradual steps over timescales of minutes... Ganesh found problem linked to gain changes caused by bias cable movement, fixed. Tsys stable now at 25.9K/27.9K for A/B.
ansanolla@naic.eduMonday, February 21, 2005 15:43:45Coco found that system was warming up , so he hooked up the vacuum pump on Dewar.
Bill Genterwgenter@naic.eduSaturday, July 2, 2005 08:15:16Yesterday afternoon the SBN receiver warmed up following a power dip. It appears that the vacuum was lost at this time. Ganesh connected the vacuum pump at about 6:15 PM. This morning I arrived at 6 AM to work on the receiver, but due to an experiment in progress and blocking, I was unable to get to the receiver priot to 7:20 AM. The vacuum was at 2uC and I started the cool down. The receiver should be ready for use by noon today.
W. Genterwgenter@naic.eduMonday, July 18, 2005 14:16:19This morning during system checks we found that the Dewar was warming up. The vacuum had again failed. Placed the turbo vacuum pump on the Dewar. The system is cooling once again.
W. Genterwgenter@naic.eduFriday, July 22, 2005 15:35:11Checked new converter for Jaun-like. All OK
ansanolla@naic.eduFriday, August 26, 2005 13:05:23Due to power failure on 8/25/2005 system warmed up. Been cooled.
ansanolla@naic.eduTuesday, August 30, 2005 16:00:10System back to normal.
W. Genterwgenter@naic.eduWednesday, September 7, 2005 09:51:19Four Lakeshore Temperature diodes have arrived for the new SBN Receiver.
ansanolla@naic.eduThursday, October 20, 2005 13:09:20Dewar lost vacuum,so receiver warmed up.Was pumped and it returned to normal operation.
ansanolla@naic.eduFriday, October 21, 2005 11:09:33Pump still on it until Monday.
ansanolla@naic.eduMonday, October 31, 2005 14:57:53The sys on channel A was 29k instead 25k.JRV open the wg to check humidity at the Dome,but it was dry.So after doing this T sys was checked again and then it returned to 25k , what it was ,we don't know.
ansanolla@naic.eduFriday, November 4, 2005 11:28:12System cryohead turn off, system warming up.Going to be cooled on Monday 7 ,2005.
ansanolla@naic.eduMonday, November 7, 2005 14:35:55SBN cooled down by WG. Vacuum valve closed.
ansanolla@naic.eduTuesday, November 15, 2005 09:21:02Dewar bias cable replaced. Former one had 2 open wires (21 and 34) and 1 intermmitant (11).
W. Genterwgenter@naic.eduWednesday, December 14, 2005 09:04:1730MHz baseband mixer in rack 8-10 found not to have output.This delayed this mornings run.
ansanolla@naic.eduMonday, March 13, 2006 14:17:16System down due to modification of post amp lay out.Now post amp will have the abilisy to choose between an fo of 2270 mhz or 2380 Mhz.In other way a filter bank was added.It was a NASA request to add the 2270 Mhz to the system.
ansanolla@naic.eduWednesday, March 15, 2006 11:44:06System back ,now you can choose between 2290 Mhz or 2380 Mhz operation , the Fo swap have to be performed upstairs at the post amp chassis.Both frecuencies were calibrated.
ansanolla@naic.eduThursday, March 16, 2006 14:37:15The channel A rf connector at the output of the post amp was loose. Tigh it up.
ansanolla@naic.eduTuesday, May 9, 2006 14:19:25Channel B of SBN recvr connected to the SN xmitter new horn in order to check its Tsys.
ansanolla@naic.eduThursday, July 6, 2006 16:20:09The recvr Dewar lost its vacuum , been pumped.Vacuum down to 7 un's after 5 minutes of pumping.
ansanolla@naic.eduFriday, July 7, 2006 11:21:13Dewar vacuum below 1 un.Dewar vac valve closed.An eye on it.
W. Genterwgenter@naic.eduMonday, July 24, 2006 10:52:41This receiver was repoted warm by the Telescope Operator at 0400 this date. On investigation we found that the Dewar had lost vacuum. The turbo pump was applied and in less than 10 minutes we had a vacuum of <2uc. We will recool the receiver on 7/26/06.
ansanolla@naic.eduWednesday, July 26, 2006 11:31:13Receiver been cooled down.
ansanolla@naic.eduWednesday, July 26, 2006 11:35:25Forgot to mention that a new film of vacuum seal was added to the leak in the vacuum valve.
ansanolla@naic.eduThursday, July 27, 2006 16:31:28System up again.
W Genterwgenter@naic.eduTuesday, August 22, 2006 10:02:04Saturday, August 19, 2006 M. Nolen placed the receiver center frequency at 2290MHz. Phil changed the Tsys settings to reflect this change. Monday, August 21, 2006 in the AM the receiver center frequency was changed to 2380MHz. Today Phil changed the Tsys to reflect the change.
ansanolla@naic.eduWednesday, November 15, 2006 15:56:14System out of service , the cryo head is locked up. We tried this morning a purge due to contamination on it.After the purge we started a cool down , but in the afternoon we found the crosshead locked.Need replacement.
ansanolla@naic.eduThursday, November 16, 2006 13:03:18Cryohead locked up was replaced with the spare one.System being cooled.
ansanolla@naic.eduFriday, November 17, 2006 11:22:03System back to normal.Cryo head was replaced.
ansanolla@naic.eduMonday, November 27, 2006 12:57:24System cryo turned off in order to help the XB system to cool down , having problems with it. To be cooled down on Wednesday.
ansanolla@naic.eduWednesday, November 29, 2006 14:19:55System back to normal , was cooled down.
ansanolla@naic.eduTuesday, December 5, 2006 15:25:06System has being down since Saturday due to a leak in the cryo head of the XB receiver , it drained out all the He of these cryo system (XB-SBN).It is being cool down with the CBH receiver cryo comp & lines.
ansanolla@naic.eduWednesday, December 6, 2006 10:22:47Vac valve closed.System back to service.
crvcrios@naic.eduTuesday, December 12, 2006 13:24:40System now using SBH cryo Comp. and lines.(ed009, lines:17-18)
CRVWednesday, January 2, 2008 14:28:54Vacuum pump connected to dewar and pumping. Dic.5th,2007
CRVWednesday, January 2, 2008 14:30:35Dewar cool down Started Dic.6th,2007
CRVWednesday, January 2, 2008 14:31:25Dewar cold. Dic.6th,2007 afternoon
CRVWednesday, January 2, 2008 14:35:08Dewar warmed out because the cross head broke three screw. The compressor lost all the pressure and it also affected XB dewar. At the end of the Day the system was cleaned the screws replaced and SBN cooling down. 12/21/2007
crvTuesday, January 29, 2008 16:59:06The crosshead mechanicaly failed yesterday 1/28/08
crvTuesday, January 29, 2008 17:01:08the system was vac. crosshead replaced and cooling. 1/29/08
crvWednesday, February 6, 2008 14:06:10Reciver was changed with reciever sbl. The system was originally with compressor ED008 and is now with compressor 009.
Dana Whitlowdwhitlow@naic.eduThursday, February 14, 2008 14:34:37Swapped crosshead driver with that of SB, on 13 Feb 2008. SBN crosshead was in distress, making lots of humming noises and sometimes chattering and freezing.
CRVWednesday, November 5, 2008 13:11:41on monday Nov. the 3erd the system was racheting like it was contaminated. The system was warmed because the purging method used during last week didn't work.In the afternoon Guarionex and I removed the cryo pump to verify the contamination. The dewar and the cryo pump were cleaned and the system started cooling at 5pm.
CRVWednesday, November 5, 2008 13:13:23System vacuum valve was closed and the pump was removed. The system temp was 21K before the monday and is in 13.4K today.
crvWednesday, July 1, 2009 14:39:04Cryo pump was replaced due to contamination. system was cleaned ,vacuumed , charge and cooled down on 5/14/2006
crvcrios@naic.eduMonday, August 10, 2009 10:14:06dewar was vca. today and started cooldown when vac. wAS ON 2 microns.
crvrios@naic.eduMonday, August 10, 2009 14:16:24Vac. pump removed from dewar cryo temp. at 15.8 K.
crvMonday, September 27, 2010 17:03:11system was warmed and purged an recooled is operational.
criosMonday, May 23, 2011 16:16:36The system was sounding contaminated the dewar was partially warmed up and purged. The system is already cold.
Carlos Rioscrios@naic.eduThursday, February 2, 2012 10:23:36The receiver cooling system was swaped with the XB. Connecting the XB with the 327Mhz receiver. SBN is now alone with the small compressor.This was the SBN original setup. January 26th 2012
crioscrios@naic.eduMonday, May 7, 2012 10:16:38The dewar had been warming up due to bad vacuum. We turned the cooling off on friday May 4th 2012. The receiver is cooling now.
criosFriday, August 10, 2012 17:06:56Post amp/filter bank chasis reinstalled. Was been used for the last 3 days to test the new S-Band receiver.
Denis urbaindurbain@naic.eduTuesday, September 11, 2012 08:09:12Cold head was contaminated. The receiver was warmed up and the coldhead removed. Traces of water were found inside the canister. The displacer was cleaned and the canister was honed and cleaned. The vacuum was restored and the receiver was cooled again.
CRIOSFriday, September 21, 2012 10:44:26shutter removed to be installed and tested on new SBN
Denisdurbain@naic.eduFriday, October 5, 2012 11:24:43The cold head was again contaminated, traces of water were found when we removed the displacer.The displacer was baked for a few hours and the cold head was reassembled and flushed.
Denisdurbain@naic.eduFriday, October 5, 2012 11:34:00The shutter controller box and motor had been removed to test the new S Band narrow receiver shutter mechanism. After completion of the test, the controller was reinstalled with new limit switches and the tracks were cleaned. The cables going from the control box to the limite switches have been replaced as well.
Denis Urbaindurbain@naic.eduWednesday, January 23, 2013 16:25:59The new S Band Narrow receiver was installed at the end of October 2012. After the initial setup, the receiver was moved up 1.2in to improve the focal position. Some sign of contamination in the refrigerator forced us to warm-up the receiver last week. When the displacer was removed, traces of water were found on the displacer. The displacer was cleaned, baked and reinstalled, while we were pumping on the Dewar. After the cool down the displacer was knocking, to prevent any damage to the cold head, the system was turned off and warmed-up again. Guarionex checked the tolerances and turned the displacer 180deg before installing the cold head again.
Denis Urbaindurbain@naic.eduMonday, February 25, 2013 13:17:04The coldhead got contaminated just before an S band run. We tried to minimize the down time by swapping the coldhead with the one from the test Dewar. Water was found on the displacer. We installed the newly build de-hydrating cartridge from NRAO and swap compressor to avoid more problem with the receiver. Unfortunately the coldhead got contaminated again and the receiver was warmed up again during the weekend. The coldhead after the second flush was still making noise and the compressor was swapped with the compressor from the S band wide because the return pressure was very low and might have been affecting the refrigerator. We managed to keep the coldhead running for the duration of the S band run (one week) before it was turned off and left with the vacuum pump connected over the weekend. The coldhead was replaced and the vacuum being good at 2 millitorr, the cool down was started immediately. The coldhead was running smoothly with no sign of knocking.
Denis Urbaindurbain@naic.eduFriday, August 9, 2013 16:57:40The shutter was giving some problems, the movement was jerky. The limit switches were replaced with no result, further investigation pointed to the cable feed-thrus on the blue cabinet. The feed-thrus were replaced and after that the shutter was working fine. After a S band radar experiment both amplifiers got damaged, the first stage on both amplifier are dead. We have to investigate if the shutter could have given wrong information and the Transmitter could have been on while the receiver was not covered.
Denis Urbaindurbain@naic.eduWednesday, September 25, 2013 09:32:07We think we found the cause of the amplifier failure. During AO19 campaign, we tried to run a Radar experiment but we had to by-pass some of the interlock controls because of the missing receivers. We disabled by mistake the closed position indicator for the shutter. A problem with the transmitter made us lose the synchronization between the rotation of the floor, the transmitter and the shutter and we apparently did transmit while the shutter was not fully closed.
RFI and Post amp chasis removalcrios@naic.eduThursday, March 20, 2014 15:27:05Phil was searching for a RFI in the dome and found a posible cause on the postamp chassis. Chassis brought down to be tested on the lab.
Cold-head contaminationfelix.fernandez@naic.eduTuesday, April 22, 2014 17:19:16Today,since the compressor driving the SBN, SBL and CBH was leaking He, it was decided to move the SBN and SBL receivers to the M600 compressor that typically is used for ALFA (note: ALFA is currently down for repairs). After 15-20 minutes of operation the SBN and SBL temperatures started rising. Once we got to the G. Dome we found both cold-heads to be severely contaminated. Thus, a thermal cycle was started today afternoon at 15:30.
SBN back onlinefelix.fernandez@naic.eduFriday, April 25, 2014 06:54:06SBN was back to operational status on 4-24-2014.
Replaced post-amp chassisfelix.fernandez@naic.eduThursday, May 8, 2014 14:47:00The SBN had some RFI issues coming from its post-amp chassis, specifically the noise injection pcb (PCB: PC950622). There was an HCT04 Hex-inverter chip (U2) that was not been used and its outputs left floating. Furthermore, the chip itself was damaged and had severe oscillations. The faulty chip has been removed and the chassis was replaced.
Amplifier replacementfelix.fernandez@naic.eduMonday, July 7, 2014 16:35:09During this weeks window we will replace the SBN amplifiers by those that were repaired at Chalmers. These replacements are SN03 and SN02. These were repaired this last winter and have a new breed of InP transistors.
Compressor failurefelix.fernandez@naic.eduWednesday, July 30, 2014 11:09:56The SBN, SBL/W , and CBH compressor (model M600) went "dead" after last Jun-19-2014 power failure. After several attempts to fix the problem while the compressor was upstairs it was concluded that the compressor's scroll pump was seized and was tripping the over-current sensor. Hence, SBN, SBL/W and CBH have been warm since that event. Today we finally got the opportunity to replace the compressor. All the aforementioned receivers will begin their vacuum and cool-down cycles sequentially today.
Correction to July 7 log entryfelix.fernandez@naic.eduFriday, August 1, 2014 08:31:10On Monday, July 7, 2014 16:35:09 log entry it was mentioned that we replaced both amplifier in the SBN. Although this was the original plan, we only replaced one amplifier on Pol A. The amplifier on Pol A was replaced by Chalmers SN03.
Oscillating Temperature Readingsfelix.fernandez@naic.eduMonday, August 4, 2014 14:30:54There has been some recent event related to the SBN's temperature monitoring system. Mainly there are two strange behaviors: 1. the readings in the control room seem to become unstable for short period of time. The readings will be fine for several hours and the suddenly they will oscillate from absurdly low and high values. 2. when reading the Lakeshore monitor upstairs the readings will start fine but then slowly start creeping up. Today we went up and observed with the oscilloscope the signals coming from the Dewar and they seemed stable and without oscillations. We then decided to replace the PCB AO940330. This PCB multiplexes the temperature diodes' signal to the Post-Amp chassis. The readings seemed to be stable for a few hours but started to oscillate once again.
Temperature readings oscillations felix.fernandez@naic.eduTuesday, August 5, 2014 13:16:56Today we found out that the ANALOG OUTPUT connector from the bias chassis has a broken connection inside the chassis. There was no signal or intermittent signal getting to the GIP MUX at the turret floor. The connector was repaired and temperature readings oscillation seemed to have stooped. There still remains the constant creeping up of temperature readings in the Lakeshore M200. We will look further into this issue at latter time.
Pol-B RF cable's ground shield repairfelix.fernandez@naic.eduWednesday, August 6, 2014 15:31:45Today we found out that the ground shield on Pol-B's heliax RF cable was loose. The cable was striped and the ground shield was properly reattached. The repaired cable is the one connecting the post-amp chassis to the blue cabinets on the turret floor.
Bad Tsys and low gain on Pol-Bfelix.fernandez@naic.eduTuesday, October 14, 2014 11:19:56On Oct-12-2014 the gain of Pol-B appeared to be 10dB lower than normal. Also, Pol-B's Tsys reading were much higher than normal. After some debugging we found that disconnecting and reconnecting the RF cable on the Post-Amp Chassis fixed the problem. NOTE: There is some new RFI inside the band of interest.
Cold-head replacementfelix.fernandez@naic.eduTuesday, February 10, 2015 11:49:45SBN's temperature had been increasing for the past few days. Yesterday the 16K stage was at 35K. It was deemed necessary to warm-up the receiver to service the cold-head. The receiver was warmed-up yesterday and today we replaced the cold-head. NOTE: We used the new bench-top Varian pump and the vacuum progressed from atmosphere to 10 mTorr within 2 hours. This will allow us to start the cool-down process much earlier than expected.
SBN back onlinefelix.fernandez@naic.eduWednesday, February 11, 2015 15:01:02SBN is back to operational status after 8hrs of vacuum using the TPS-bench pump and another 24hrs of cool-down. The only work done was replacing the cold-head.
5dB noise florr difference when the sb750nb is setup is selectedfelix.fernandez@naic.eduFriday, May 15, 2015 10:53:03M. Nolas informed that there was about 5dB difference in Pol-B's gain during MAY-13 radar run. We found out that there was indeed a 5dB gain drop in Pol-B after the second post-amp chassis. This 5dB drop was due to having 20dB pad on Pol-B's input and a 15dB pad on Pol-A's input. We then installed a 15dB pad on Pol-B's input. Furthermore, we noticed a that Pol-B's noise floor was about 5dB higher that Pol-A's. This delta is only present when using the 750MHz channel without the narrow-band option. We will look further into this situation early next Monday.
Cold-head replacementfelix.fernandez@naic.eduTuesday, August 11, 2015 07:40:01On Aug-9-2015 at 3PM the SBN suddenly started to warm up. The cold-head was replaced on Aug-10-2015. After replacing the cold-head it was concluded that the seals on the previous cold-head had worn out.
Cold-head contaminationfelix.fernandez@naic.eduFriday, January 8, 2016 09:21:25Before the start of the WD campaign on JAN-4 the SBN cold-head started ratcheting. We decided to warm up the receiver during the first part of the WD campaign and today (JAN-8) we replaced the cold-head and started the cool-down.
Lost bias/physical temp monitoring felix.fernandez@naic.eduThursday, March 30, 2017 11:39:25During MAR/29 and MAR/30 there was no bias/power/physical temp. monitoring. All values reported during that time period were junk. On MAR/30 a cable was found to have become loose. After reconnecting the cable all values came back to their normal range. The cable probably came loose during servicing of the 327 Rx.
Intermittent shutter "OPEN" behaviorfelix.fernandez@naic.eduTuesday, May 9, 2017 08:25:38For the past few months every so often the Tsys values would report unlikely low values (i.e. 12K). This became more persistent during the past two weeks. Phil figured out that the shutter system was failing on the "OPEN" command. C. Rios and R. Alicea checked the system for possible causes and found a faulty relay (model: W88ANCPX-2) which had intermittent connections inside. The relay was replaced on May 8, 2017.
Cool-down started`felix.fernandez@naic.eduWednesday, October 18, 2017 08:07:19Started the SBN cool-down after hurricane Maria.
Col-head Ratcheting felix.fernandez@naic.eduThursday, October 19, 2017 12:55:21The Rx's cold-head was ratcheting during today's inspection. The system will be purged in an effort to resolve the issue.
Sudden gain drop on PolBfelix.fernandez@naic.eduMonday, July 23, 2018 10:31:19SBN event 180718 * Got a call from control roonm on 180717 at 5:36PM. The operator (E. Mu~iz) informed me that there was an issue with the SBN Rx. Since the observer (S. Marshall) was there I asked to speak with him. S. Marshall mentioned that they had been dealing with low signal returns from their observations. As a debug trial he thinks he might have opened the RX shutter while still transmitting. I think this is impossible due to the interlock system but it has happened before (late 2013 both SBN LNA were damaged because the interlock system had been bypassed during an AO19 camapaing). * I went to AO and started to debug the situation. A. Tsys SBN polA = 25K, polB = 80K B. PolB bandpass at Spectrum Analyzer (SA) was about 10dB below PolA. Also the strong RFI located at the high side edge was lower. C. Swapped the signals at the Dome XFER switch (zmon switch?)and the problem shifter to PolA. D. Took the Anritsu SA up to the Dome and measured at different points i. Filter Chassis' ouput of the showed the same 10dB drop of PolB ii. Dewar's output had a similar drop * Concluded thar the problem resides within the Dewar E. Asked the operator to warm-up the Rx at midnight (new operator was D. Padilla) * All control room measurements were done with ifsetup sb750nb command * On 180718 wen to the Dome A. Tsys sbn polA = 107K, polB = 95K B. SBN showed the same passband response on both channels (yet the RFI was smaller by about 5dB on PolB), using the control rooms SA C. Used the R&Z network analyzer (NA) to test the Rx i. Filter chasis showed about the same gain and passband response. There was a 1.7dB gain difference and some minor differences at the passband's low side ii. Post Amp chassi, exactly the same gain and passband iii. Filter + Post Amp chassis, had the same gain and freq response. iv. Dewar's output were about the same level * It seems the increase in physical temp leveled the LNAs performance. Could this be a cyogenic related issue? Opted no to replace the LNA yet. We will do a thermal cycle hoping this will revert PolB's noise temp to typical values. Started cooldown at 5:50PM on 180718 with the vaccumm reading of 10mTorr 180719 at 7:50AM SBN physical temp was at 87K a Tsys reported back both polarizations at 55K. Both pol frequency responses using ifsetup sb750nb were as expected. Also we did an adjpwrif1 and adjpwrif2 and the system did balance out as expected. So far it seems the system is healty. The cooldown rate is about 10K drop per hour. Within the next 8hrs the system will reach its minimum temp. 180719 at 2:30PM SBN physical temp 27K, Tsys showed around 33K on each pol. System progresses as expected. We hit the temp knee at about 2PM, usually on SBN (due to the temp diode offset reading) it takes about 8hr to reach min temp (there will be another inflexion and suddenly drop to about 8.9K) We should do another Tsys at about 10PM. 1180719 at 10PM all Tsys reported to be 24K. We are good to go.
Thermal Cyclefelix.fernandez@naic.eduWednesday, January 16, 2019 10:22:30During 12-jan-19 there was a power dip and an unexpected thermal cycle started on SBN,SBW,LBW. Seems to be compressor related.
Thermal Cyclefelix.fernandez@naic.eduWednesday, April 17, 2019 09:06:12Another set of power dips caused a loss of vaccumm. The systems started to warm up at 01:30 on 2019-04-17. We started to evacuate the system at 08:00 and the cool down should start sometime today.
Heater installationfelix.fernandez@naic.eduWednesday, April 24, 2019 13:03:15After the last unexpected thermal cycle we opted to do a controlled one. This new thermal cycle would include a evacuating the dewar as we heat up the inside. Usually all of our dewars have internal heaters for such purpose. Today we found out that SBN has no internal heater, so we will install one today. Then we will start the system evacuation.
Controlled thermal cyclefelix.fernandez@naic.eduWednesday, May 8, 2019 11:40:33After 2019-05-04 observation we did a controlled thermal cycle to replace the system's cold-head. The cold-head was replaced on 2019-05-06 and the nominal physical temperature of "10K" (there is an offset in the temperature reading) was reached on 2019-05-07.
PolB Tsys is 10K higherfelix.fernandez@naic.eduWednesday, May 8, 2019 11:42:47The last Tsys shows SBN's polB about 10K higher that polA. We looked into this and found out that there is about a 10dB difference in SNR from polA to polB, polB having a lower SNR. We went up to the dome and found that the dewar's outputs (i.e. LNA outpus) had the same 10dB SNR difference. Hence, something happened during the last thermal cycle and the LNA's performance are no longer symmetrical. We will address this as soon as possible.
Note about Pols in SBNfelix.fernandez@naic.eduFriday, May 10, 2019 11:23:48The SBN has its pols swapped somewhere down the IF chain. That is, PolA from the post amp chassis shows up as PolB in our downstairs IF/LO system. Hence, the current issue with a degraded SNR, which shows up in PolB at the downstairs IF/LO system is actually coming from PolA at the post amp chassis. Also, PolB is CH2 in our "planetary down-converter", and this is the preferred polarization during SB radar observations.
LNA issues continuefelix.fernandez@naic.eduWednesday, March 4, 2020 14:06:41The problem with SBN's LNAs continue. We ended bringing down the receiver for a detailed inspection. All passive elements proved to be within specs. The problem is an LNA which has lost some performance. This LNA currently is amplifying the same-circular (SC) signal and it is being sourced from CH2 of the "Planetary Downconverter". We ultimately need t o purchase new LNAs since the spare LNA (which we thought was dead) is showing the same performance issues.

Back to Top